# The Economics of Blockchain Governance: Evaluate Liquid Democracy on the Internet Computer

### v1

@@ Abstract
Decentralized Autonomous Governance (DAO), Internet Computer Protocol (ICP), Liquid Democracy, AI Governance, Cooperative AI, Blockchain
@@ 4 Conclusion and Future Research
  • Market Sentiment and Behavioral Analysis: Integrating market sentiment analysis [13, 14, 15] could offer insights into investors’ perceptions and emotional responses to token developments. This understanding could deepen the analysis of how market sentiment directly influences staking decisions and contribute to a nuanced interpretation of staking behaviors.
  • User Preferences and Risk Tolerance: Employing behavioral economics methodologies, such as surveys or experiments, to assess user preferences and risk tolerances [16, 17] could yield valuable insights. This approach could reveal how different risk profiles among stakeholders within the Web3 ecosystem affect staking ratios under varying conditions of inflation rates, age bonuses, and dissolve delays.

### v2

@@ Abstract
Decentralized Autonomous Governance (DAO), Internet Computer Protocol (ICP), Liquid Democracy, Agent-based Modeling, Cooperation, Blockchain
@@ 4 Conclusion and Future Research
  • Market Sentiment and Behavioral Analysis: Integrating market sentiment analysis [13, 14, 15, 16] could offer insights into investors’ perceptions and emotional responses to token developments. This understanding could deepen the analysis of how market sentiment directly influences staking decisions and contribute to a nuanced interpretation of staking behaviors.
  • User Preferences and Risk Tolerance: Employing behavioral economics methodologies, such as surveys or experiments, to assess user preferences and risk tolerances [17, 18, 19] could yield valuable insights. This approach could reveal how different risk profiles among stakeholders within the Web3 ecosystem affect staking ratios under varying conditions of inflation rates, age bonuses, and dissolve delays.